OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)?

OpenStudy (anonymous):

\[r^{2} = (x-h)^{2} + (y-k)^{2}\] correct?

OpenStudy (anonymous):

no

OpenStudy (muzzack):

Standard form equation for a circle with center (h,k) and radius r is: \(\large(x-h)^{2}+(y-k)^{2}=r^2\) So you have the center, now you just need the radius. Well, given the center and a point on the circle, the radius is just the distance between those 2 points. so use the distance formula to find the distance between those 2 points, which is your r